Σ (ax, + by, + c) O B. Ос. D. ax by nxc
Consider the following linear transformation of a random variable where ux is the mean of x and o, is the standard deviation. Then the expected value and the standard deviation of Y are given as: and σχ. O B. 0 and 1 O c. cannot be computed because Y is not a linear function ofx. O D. 1 and 1
Answer #1

sum_{i=1}^{n}left ( ax_{i}+by_{i}+c ight )

に! に!

= asum_{i=1}^{n}left ( x_{i} ight )+bsum_{i=1}^{n}left ( y_{i} ight )+csum_{i=1}^{n}left ( 1 ight )

L (.ti) + b i= 1 (yi) + n × c i= 1 B

Expected value is 0

and standard deviation is 1

0 and 1 B